OpenStudy (anonymous):

the derivative of g(x)= f(x)^3 using the definition

OpenStudy (anonymous):

Are you familiar with the chain rule?

OpenStudy (anonymous):

can't use the chain rule or I would it has to be done with the definition

OpenStudy (anonymous):

Use the chain rule

OpenStudy (anonymous):

what is the derivative of x^3?

OpenStudy (anonymous):

Just to make sure, this is still the derivative of \[\left[f(x)\right]^3, \text{ right?}\] Or, is it \[f\left(x^3\right)?\]

OpenStudy (anonymous):

the way it is set up is \[f(x)^3 \]

OpenStudy (anonymous):

Okay, then that's the first one I listed. \[g'(x) = \lim_{h\to 0}\frac{\left[f(x+h)\right]^3-\left[f(x)\right]^3}{h}\] Just a moment, let me work this out on paper (if I can).

OpenStudy (anonymous):

lol ok sounds good

OpenStudy (anonymous):

is the question asking what the derivative of g(x) is?

OpenStudy (anonymous):

the derivative of \[g(x)=f(x)^3\]

OpenStudy (anonymous):

so g'(x) = f(x)^3?

OpenStudy (anonymous):

Okay, I think I've got it. However, it's not as intuitive as it should be. I'm going to ignore the limit for a while and just rewrite/rearrange the numerator for a bit (hope it doesn't get cut off): \[f(x+h)^3-f(x)^3\\ f(x+h)^3-f(x)^3 + \left[3f(x)f(x+h)^2-3f(x)^2f(x+h)\right]-\left[3f(x)f(x+h)^2-3f(x)^2f(x+h)\right]\\ f(x+h)^3-f(x)^3-\left[3f(x)f(x+h)^2-3f(x)^2f(x+h)\right] + \left[3f(x)f(x+h)^2-3f(x)^2f(x+h)\right]\\ \left[f(x+h)^3-f(x)^3-3f(x)f(x+h)^2+3f(x)^2f(x+h)\right] + \left[3f(x)f(x+h)^2-3f(x)^2f(x+h)\right]\\ \left[f(x+h)^3-3f(x)f(x+h)^2+3f(x)^2f(x+h)-f(x)^3\right] + \left[3f(x)f(x+h)^2-3f(x)^2f(x+h)\right]\\ \left[f(x+h)-f(x)\right]^3 + \left[3f(x)f(x+h)^2-3f(x)^2f(x+h)\right]\] Is that all clear?

OpenStudy (anonymous):

Darn, it did get cut off... The cut-off portion is \[...-3f(x)^2f(x+h)\] I only rewrite the first big sum, and the cut-off section reappears perfectly in the last line.

OpenStudy (anonymous):

I know it seems like I pulled that [3f(x)²f(x+h) - 3f(x)f(x+h)^2] term out of nowhere, but it's all for a reason.

OpenStudy (anonymous):

ok thats fine... I just haven't been able to figure them out

OpenStudy (anonymous):

So does it all make at least some sense now?

OpenStudy (anonymous):

umm kind of

OpenStudy (anonymous):

Okay, so the last line that I typed out is now the numerator of the limit: \[\lim_{h\to 0}\frac{\left[f(x+h)-f(x)\right]^3 + \left[3f(x)f(x+h)^2-3f(x)^2f(x+h)\right]}{h}\\ \lim_{h\to 0}\frac{\left[f(x+h)-f(x)\right]^3}{h} + \lim_{h\to 0}\frac{\left[3f(x)f(x+h)^2-3f(x)^2f(x+h)\right]}{h}\] First, let's deal with the first limit: \[\lim_{h\to 0}\frac{\left[f(x+h)-f(x)\right]^3}{h}\\ \lim_{h\to 0}\frac{\left[f(x+h)-f(x)\right]\left[f(x+h)-f(x)\right]^2}{h}\\ \lim_{h\to 0}\left[f(x+h)-f(x)\right]\frac{\left[f(x+h)-f(x)\right]^2}{h}\\ \lim_{h\to 0}\left[f(x+h)-f(x)\right]\cdot\lim_{h\to 0}\frac{\left[f(x+h)-f(x)\right]^2}{h}\\ \left[f(x)-f(x)\right]\cdot\lim_{h\to 0}\frac{\left[f(x+h)-f(x)\right]^2}{h}\\ 0\cdot\lim_{h\to 0}\frac{\left[f(x+h)-f(x)\right]^2}{h}\\ 0\]

OpenStudy (anonymous):

Now the second limit: \[\lim_{h\to 0}\frac{3f(x)f(x+h)^2-3f(x)^2f(x+h)}{h}\\ \lim_{h\to 0}\frac{3f(x)f(x+h)\left[f(x+h)-f(x)\right]}{h}\\ \lim_{h\to 0}\left[3f(x)f(x+h)\right]\cdot\lim_{h\to 0}\frac{f(x+h)-f(x)}{h}\\ 3f(x)f(x)\cdot f'(x)\\ 3f(x)^2\cdot f'(x)\] Finally done!

OpenStudy (anonymous):

ok cool thank you